OpenStudy (anonymous):

jarod is having a problem with rabbits getting into his vegetable garden, so he decides to fence it in. the length of the garden is 3 feet more than 6 times the width. he needs 132 feet of fencing to do the job. find the length and width of the garden.

OpenStudy (anonymous):

9ft

OpenStudy (anonymous):

2(6x+3)+2x=132 14x+6=132 x=126/14 x=9

OpenStudy (anonymous):

so w = the width. Since the length is 3 more than 6 times the width, you get length=6w+3. Just say length = l and you get l=6w+3. Now the perimeter is l+l+w+w or 2l+2w and equals 132 feet. Plug in 6w+3 for l and w for w and you get 2(6x+3)+2w=132

OpenStudy (anonymous):

sorry 2(6w+3)+2w=132 12w+6+2w=132 14w=126 w=9 width is 9 length is 57